One-third of the sum of the second and third consecutive numbers is 23 . What are the integers?
Answer by Theo(13342)

x is the first number
x+1 is the second number
x+2 is the third number.
1/3 * (x+1+x+2) = 23
simplify to get:
1/3 * (2x+3) = 23
multiply both sides of the equation by 3 to get:
2x+3 = 69
subtract 3 from both sides of the equation to get:
2x = 66
divide both sides of the eqution by 2 to get:
x = 33
first number is 33
second number is 34
third number is 35
34 + 35 = 69
1/3 of 69 = 23
those are your number.